# Chemical Signatures of AGB Mass Transfer in Gaia White Dwarf Companions

### v1

@@ 3.3 HIRES
We note that a subset of the MIKE and HIRES observations were originally made as part of a follow-up program of NS + MS binary candidates (K. El-Badry et al., 2024b, c), but were primarily used to measure radial velocities (RVs) rather than abundances.
3.4 Continuum normalization and RV correction
@@
We processed all reduced spectra using the iSpec package (S. Blanco-Cuaresma et al., 2014; S. Blanco-Cuaresma, 2019). To estimate the continuum for each echelle order, iSpec first applies a median and maximum filter to the data to mitigate noise effects and avoid strong absorption features, and then fits multiple splines of a specified degree to the data in each order to be used for the normalization. We chose typical window sizes of 0.050.05\,nm and 1.01.0\,nm for the median and maximum filters, respectively, and splines of degrees 4 and 5, but varied the exact choices for each order and instrument in order to optimize the continuum normalization and avoid overfitting  [^1] For more details of this process, consult the iSpec manual: www.blancocuaresma.com/s/iSpec/manual/introduction. We then merged all orders of the normalized spectra, taking an inverse-variance weighted average of the flux in overlap regions. The RV of each spectrum was calculated with iSpec’s cross-correlation algorithm using a template solar spectrum. After shifting the spectra to rest frame, if repeated observations were taken for one target with the same instrument, we co-added them into one spectrum to increase SNR.
4 Analysis
@@
Figure 2: Comparison of best-fit stellar parameters obtained in this work (Section 4.1) with those calculated using Gaia low-resolution spectra (R. Andrae et al. 2023; top row), as well as those from GALAH DR4 (S. Buder et al. 2025; bottom row). Systems with secondary masses above and below the Chandrasekhar mass of $1.4\,M_{\odot}$ are plotted with red unfilled and blue filled markers respectively.
4.1 Atmospheric parameters with excitation/ionization balance
@@
To measure equivalent widths (EWs) of absorption lines, we used REvIEW [^2] https://github.com/madeleine-mckenzie/REvIEW (M. McKenzie et al., 2022). Provided a spectrum and linelist, REvIEW fits a region $\pm 0.6\,$Å around the central wavelength of each line with up to three Gaussians to separate out contributions from closely spaced lines that may be partly blended.
@@
We then use q2333https://github.com/astroChasqui/q2 (I. Ramírez et al., 2014) to determine stellar parameters. This code is a python interface for MOOG (2019 release; C. A. Sneden 1973; C. Sneden et al. 2012) and calls the abfind function, which uses the excitation/ionization equilibrium method with the measured EWs of iron lines. This method determines the optimal set of effective temperature ($T_{\rm eff}$), surface gravity (log(g)(g)), metallicity ([Fe/H]), and microturbulence ($\xi$) for which the inferred absolute iron abundance (AA(Fe))) from individual lines is independent of the lines’ ionization states, excitation potentials (EP), and relative line strengths (i.e. reduced equivalent width, REW) (e.g. D. F. Gray, 2008).
@@
We remove lines with EW above 120120\,mÅ to avoid saturated lines, as well as those with EW below 1010\,mÅ, which we found to be dominated by noise. In cases with low SNR, we increased this lower bound to 2020\,mÅ. We use the MARCS atmosphere grid (B. Gustafsson et al., 2008) and by default, we fit all four stellar parameters. For the initial guess of $T_{\rm eff}$, log(g)(g), and [Fe/H], we used best-fit values obtained using the Gaia XP very low-resolution spectra, as described in R. Andrae et al. (2023). However, we do not use published results from R. Andrae et al. (2023) as they make use of parallaxes from the gaia_source catalog, which are derived from a single-star astrometric solution. Instead, we use re-calculated values using parallaxes from the nss_two_body_orbit table. For $\xi$, we use the empirical relation from I. Ramírez et al. (2013) (their Equation 5), derived for FGK MS stars.
@@
The first two issues are problematic as solvers are reaching the edge of the stellar atmosphere grid, making the results potentially unreliable. In addition, based on the position of our objects on the color-magnitude diagram (CMD), all primaries are expected to be MS stars with log(g)=45(g)=4-5. These issues may be traced back to inaccurate EW measurements due to low-SNR spectra, spectra broadened by stellar rotation, and/or poor choice of parameter initialization. Moreover, some stars have few strong Fe II lines with reliable EWs, which can prevent a robust constraint on log(g)(g). Additionally, $\xi$ is known to be sensitive to various systematics – such as the SNR of the spectra, continuum placement, and line blending (e.g. A. Mucciarelli, 2011; P. Jofré et al., 2014; S. Blanco-Cuaresma, 2019).
@@
In these cases, we fix $\xi$ and/or log(g)(g) to their initial values and rerun the fit, leaving the remaining parameters free. The majority of objects for which fixing log(g)(g) was required were observed by FEROS, suggesting that the issue stems from their lower SNR compared to observations made with other instruments (Section 3).
@@
In total, we obtain reliable stellar parameters for 160 targets. The remaining 18 targets have spectra that did not achieve convergence or resulted in qualitatively poor excitation/ionization balance. Some of these were obtained in poor conditions, resulting in abnormally low SNR, prohibiting accurate EW measurements. This is particularly problematic for metal-poor stars with few strong iron lines. A few stars also have broad lines indicative of rapid rotation, which can cause blending and erroneous EWs. The rest of our analysis is focused on the 160 sources with converged and reliable parameters.
4.2 Spectral fitting results
@@
In the upper panels of Figure 2, we compare stellar parameters from our analysis to those calculated with the Gaia XP spectra using the method from R. Andrae et al. (2023). We find general agreement between the two works. However, there are two notable discrepancies. Below $\sim 5500\,$K, we find a roughly linearly increasing discrepancy in $T_{\rm eff}$, where our estimates are consistently hotter. This is not unexpected given previous work which has also found that stellar parameters determined using different methods and codes tend to become more discrepant with each other at low temperatures (e.g. V. Hegedűs et al., 2023). We use the spectroscopic $T_{\rm eff}$ measurements throughout for consistency. In Appendix A.2 (Figure 16), we show that our conclusions about which sources have s-process enhancement are not sensitive to this choice.
@@
In addition, our metallicities are systematically lower by 0.2 dex on average than those estimated by R. Andrae et al. (2023). One factor that may contribute to this is that R. Andrae et al. (2023) provides the mean metallicity, [M/H], not [Fe/H], meaning that it is not a one-to-one comparison. In particular, old stars deficient in iron ([Fe/H]<1.0<-1.0) are usually more enriched in other metals, such as $\alpha$-elements (e.g. R. Cayrel et al., 2004; V. Z. Adibekyan et al., 2012) and carbon (e.g. H. Li et al., 2022), which would increase the average metal content.
@@
For objects found in the fourth data release of the Galactic Archaeology with HERMES (GALAH) Survey (S. Buder et al., 2025), we also compare the quoted stellar parameters from this catalog (with quality flags flag_sp = 0 and flag_fe_h = 0) to our results in the lower panels. We once again note an offset in [Fe/H].
@@
While these comparisons are helpful for validation, we emphasize that R. Andrae et al. (2023), GALAH, and our work all use distinct methods to estimate stellar parameters and metallicities (data-driven machine learning, synthetic spectral fitting with neural networks, and excitation/ionization balance, respectively) and different underlying radiative transfer codes, meaning systematic offsets are not unexpected (e.g. S. Blanco-Cuaresma, 2019).
@@
In Figure 3, we compare regions of the observed spectra for three objects to synthentic spectra generated using our best-fit $T_{\rm eff}$, log(g)(g), and [Fe/H]. We generated the synthetic spectra using the iSpec (S. Blanco-Cuaresma et al., 2014; S. Blanco-Cuaresma, 2019), which provides a python interface through which we employed the SPECTRUM radiative transfer code [^4] https://www.appstate.edu/ grayro/spectrum/spectrum.html (R. O. Gray & C. J. Corbally, 1994). The projected rotational velocity was fixed to the solar value ($v\sin(i)=1.6\,$km/s; e.g. Y. V. Pavlenko et al. 2012) and the resolution was set to 50,000. We did not vary individual elemental abundances when generating these models. We include these plots for objects observed with different instruments and at a range of metallicities to provide visual confirmation that the best-fit parameters obtained through our method described above are reasonable.
@@
Figure 3: Observed spectra of three targets (gray) compared with synthetic spectra generated using their best-fit stellar parameters (pink; Section 4.1). We select targets covering a wide range of [Fe/H] with similar $T_{\rm eff}$ and log(g)(g), observed with different instruments. On the left, we plot the region around the magnesium triplet, and on the right, we move to redder wavelengths populated with iron lines. Visually, we see good agreement between the observations and the corresponding models, suggesting that our stellar parameters are reasonable. For the most metal-poor star (top row), the C2 Swan band at 51655165\,Å is visible, which suggests carbon enhancement (Section 5.4).
4.3 Deriving abundances
@@
Together with the best-fit stellar parameters determined above, the EWs are inputted into q2, which returns absolute abundances of specified elements inferred by each line. We consider lines with EWs between 20 and 180180\,mÅ. Compared to iron, we set a stricter lower bound to avoid the weakest lines where noise is being fit; while we set a less restrictive upper bound as there are fewer strong lines. Most elements of interest have at least one line with EW below this looser upper limit, and we have confirmed that the inferred abundances from individual lines with EWs above and below 120120\,mÅ are consistent, suggesting that saturation is not a major issue. In Table 1, we list key lines of elements used to determine their abundances which are discussed in Section 5. Note that we completely exclude several strong lines (e.g. Ba II $\lambda$4554Å) as these are often saturated, making their EWs uninformative.
For each element, we discard lines which have residuals of more than 0.3 dex from the median. Unless otherwise stated, we use solar abundances from M. Asplund et al. (2021) to calculate the differential abundances. Where there are multiple lines, the best-fit value is taken to be the median of the abundances inferred from each line and the error is taken to be their standard deviation divided by the square root of the total number of lines. If only one line is measured, we set a fixed error of 0.2 dex. This is a conservative choice, being roughly the 90th percentile of the standard deviations of abundances with multiple lines.
We note that as expected, there is close agreement between [Fe I/H] and [Fe II/H] for most of our systems. However, approximately 17% of them have discrepancies between greater than 0.20.2\,dex. Almost all of these are cases where log(g)(g) was fixed to achieve convergence of stellar parameters (Section 4.1). Since log(g)(g) is determined via ionization balance, this is unsurprising.
@@
5 Results
The resulting best-fit stellar parameters and abundances for all 160 objects are summarized in Tables 2 and 3 in Appendix A.3.
5.1 Abundance patterns
@@
Panels (B) to (D) show $\alpha$ element abundances relative to iron, $\alpha$/Fe], against [Fe I/H]. Our measured abundances follow the typical anti-correlation between $\alpha$/Fe] and [Fe/H] observed in the Galactic disk (e.g. M. R. Hayden et al., 2015), with a high-$\alpha$ plateau at low [Fe/H] due to rapid enrichment of stars formed at early times by core collapse supernovae (e.g. E. Delgado Mena et al., 2017; S. Buder et al., 2021). This is one indication that our measured abundances are reasonable.
@@
In the right panel of Figure 6, we plot [Ba/Y] (i.e. [Ba/Fe] - [Y/Fe]) as a function of the metallicity, [Fe/H]. This is a measure of the relative abundance of heavy (e.g. Ba, La, Ce) to light s-process (e.g. Y, Zr) elements (often denoted as [hs/ls]), which is commonly used to compare observations to predicted yields from AGB nucleosynthetic models (e.g. M. Lugaro et al., 2003; D. M. Allen & B. Barbuy, 2006). Above [Fe/H]=1.0=-1.0, almost all stars are consistent with negative [Ba/Y] within their observational uncertainties, while the three stars with [Fe/H]<1.0<-1.0 have positive [Ba/Y]. This implies that the yields of the light s-process element, Y, is greater than heavy s-process element, Ba, in AGB donors, unless they are very metal-poor. Although this trend is generally consistent with the literature, we note that there is no clear inverse correlation between the two above [Fe/H]=1.0=-1.0, in contrast to results from previous works (e.g. R. Smiljanic et al., 2007; S. Cristallo et al., 2016; B. Cseh et al., 2018; G. Yang et al., 2024).
@@
We find no clear separation in the abundance patterns for systems in the NCE vs. RCE samples, which suggests that – within our sample, which is restricted to RCE stars with relatively small color excesses, low eccentricity, and companion masses near $0.6\,M_{\odot}$ – they probe the same underlying population. In particular, there are barium dwarfs in the RCE sample, which would not be expected if they were triples containing tight MS+MS binaries. This reflects the fact that the cut on color excess removes some fraction of true WD + MS binaries from the non-class I sample, as discussed in N. Yamaguchi et al. (2025).
@@
5.2 Dependence on orbital and stellar parameters
In Figure 8, we show the measured [Ba/Fe] against several orbital and stellar parameters: orbital period ($P_{\rm orb}$), photocenter semi-major axis ($a_{0}\times d$, where $a_{0}$ is the angular size in mas and $d$ is the distance in kpc; this is equal to the primary semi-major axis for a dark secondary), eccentricity ($e$), MS star mass ($M_{\rm 1}$), and companion mass ($M_{\rm 2,dark}$). $M_{\rm 1}$ was obtained from Gaia’s binary_masses catalog. $M_{\rm 2,dark}$ is the mass inferred from the astrometric orbit assuming a completely dark companion, which is a reasonable assumption for a cool WD or NS. To estimate the uncertainty, we use Monte Carlo sampling, generating 100 realizations of $M_{\rm 2,dark}$ by drawing from Gaussian distributions of the input parameters ($a_{0}$, $P_{\rm orb}$, $M_{\rm 1}$, and parallax $\varpi$) and taking the standard deviation of these realizations. The histogram under each plot shows the fraction of barium dwarfs (as defined in Section 5.1) over the total number of stars in our sample. We selected bins so that the number of systems in each bin is approximately constant. The same plots for [Y/Fe] can be found in Figure 17 in Appendix A.4.
Overall, we see no strong dependence of either [Ba/Fe] or the barium dwarf fraction with orbital period or semi-major axis. There is no strong trend with eccentricity at $e<0.3$, but at higher eccentricities, we find only one barium star (Section 5.3). This is likely due to highly eccentric systems hosting NS or high-mass WD secondaries, which originate from more massive progenitors with lower s-process yields (Section 7).
In panels (D) and (E.2), we see that for barium dwarfs, [Ba/Fe] increases with both increasing primary and secondary mass, consistent with the findings of P. Rekhi et al. (2024, 2025). It is important to point out that these quantities are not independent of each other. More massive primaries require more massive WDs to rule out single MS companions and enter the non-class I sample.
However, the trend with secondary mass eventually falls off for masses above $\sim 0.8\,M_{\odot}$. Moreover, across the entire range of companion masses, there are stars which show no s-process enhancement. This suggests that mass is not the sole determining factor for whether or not stars end up as barium dwarfs. We discuss possible physical explanations of the observed trends and scatter in Section 7.

### v2

@@ III.3 HIRES
We note that a subset of the MIKE and HIRES observations were originally made as part of a follow-up program of NS + MS binary candidates (El-Badry et al., 2024d, c), but were primarily used to measure radial velocities (RVs) rather than abundances.
III.4 Continuum normalization and RV correction
@@
We processed all reduced spectra using the iSpec package (Blanco-Cuaresma et al., 2014; Blanco-Cuaresma, 2019). To estimate the continuum for each echelle order, iSpec first applies a median and maximum filter to the data to mitigate noise effects and avoid strong absorption features, and then fits multiple splines of a specified degree to the data in each order to be used for the normalization. We chose typical window sizes of 0.050.05\,nm and 1.01.0\,nm for the median and maximum filters, respectively, and splines of degrees 4 and 5, but varied the exact choices for each order and instrument in order to optimize the continuum normalization and avoid overfitting  [^1] For more details of this process, consult the iSpec manual: www.blancocuaresma.com/s/iSpec/manual/introduction. We then merged all orders of the normalized spectra, taking an inverse-variance weighted average of the flux in overlap regions. The RV of each spectrum was calculated with iSpec’s cross-correlation algorithm using a template solar spectrum. We provide a table of RVs in Appendix A.1. After shifting the spectra to rest frame, if repeated observations were taken for one target with the same instrument, we co-added them into one spectrum to increase SNR.
IV Analysis
@@
Figure 2: Comparison of best-fit stellar parameters obtained in this work (Section IV.1) with those calculated using Gaia low-resolution spectra (Andrae et al. 2023; top row), as well as those from GALAH DR4 (Buder et al. 2025; bottom row). In the upper left of each panel, we note the median and standard deviation of the offsets ($\Delta$ = This work - Gaia XP/GALAH). For plots comparing [Fe/H], we color the points by the value of the offsets in $T_{\rm eff}$. For the remaining panels, we show systems with secondary masses above and below $1.4\,M_{\odot}$ with red unfilled and blue filled markers, respectively.
IV.1 Atmospheric parameters with excitation/ionization balance
@@
To measure equivalent widths (EWs) of absorption lines, we used REvIEW [^2] https://github.com/madeleine-mckenzie/REvIEW (McKenzie et al., 2022). Provided a spectrum and linelist, REvIEW fits a region $\pm 0.6\,$Å around the central wavelength of each line with up to three Gaussians to separate out contributions from closely spaced lines that may be partly blended.
@@
We then use q2333https://github.com/astroChasqui/q2 (Ramírez et al., 2014) to determine stellar parameters. This code is a python interface for MOOG (2019 release; Sneden 1973; Sneden et al. 2012) and calls the abfind function, which uses the excitation/ionization equilibrium method with the measured EWs of iron lines. This method determines the optimal set of effective temperature ($T_{\rm eff}$), surface gravity (log(g)(g)), metallicity ([Fe/H]), and microturbulence ($\xi$) for which the inferred absolute iron abundance (AA(Fe))) from individual lines is independent of the lines’ ionization states, excitation potentials (EP), and relative line strengths (i.e. reduced equivalent width, REW) (e.g. Gray, 2008). q2 calculates errors on the stellar parameters by computing the effect of small perturbations in these parameters to the equilibrium conditions, following the methods of Epstein et al. (2010) and Bensby et al. (2014).
@@
We remove lines with EW above 120120\,mÅ to avoid saturated lines, as well as those with EW below 1010\,mÅ, which we found to be dominated by noise. In cases with low SNR, we increased this lower bound to 2020\,mÅ. We use the MARCS atmosphere grid (Gustafsson et al., 2008) and by default, we fit all four stellar parameters. For the initial guess of $T_{\rm eff}$, log(g)(g), and [Fe/H], we used best-fit values obtained using the Gaia XP very low-resolution spectra, as described in Andrae et al. (2023). However, we do not use published results from Andrae et al. (2023) as they make use of parallaxes from the gaia_source catalog, which are derived from a single-star astrometric solution. Instead, we use re-calculated values using parallaxes from the nss_two_body_orbit table (Appendix A.2). For $\xi$, we use the empirical relation from Ramírez et al. (2013) (their Equation 5), derived for FGK MS stars.
@@
The first two issues are problematic as solvers are reaching the edge of the stellar atmosphere grid, making the results potentially unreliable. In addition, based on the position of our objects on the color-magnitude diagram (CMD), all primaries are expected to be MS stars with log(g)=45(g)=4-5. These issues may be traced back to inaccurate EW measurements due to low-SNR spectra, spectra broadened by stellar rotation, and/or poor choice of parameter initialization. Moreover, some stars have few strong Fe II lines with reliable EWs, which can prevent a robust constraint on log(g)(g). Additionally, $\xi$ is known to be sensitive to various systematics – such as the SNR of the spectra, continuum placement, and line blending (e.g. Mucciarelli, 2011; Jofré et al., 2014; Blanco-Cuaresma, 2019).
@@
In these cases, we fix $\xi$ and/or log(g)(g) to their initial values and rerun the fit, leaving the remaining parameters free. The majority of objects for which fixing log(g)(g) was required were observed by FEROS, suggesting that the issue stems from their lower SNR compared to observations made with other instruments (Section III).
@@
In total, we obtain reliable stellar parameters for 161 targets. The remaining 17 targets have spectra that did not achieve convergence or resulted in qualitatively poor excitation/ionization balance. Some of these were obtained in poor conditions, resulting in abnormally low SNR, prohibiting accurate EW measurements. This is particularly problematic for metal-poor stars with few strong iron lines. A few stars also have broad lines indicative of rapid rotation, which can cause blending and erroneous EWs. The rest of our analysis is focused on the 160 sources with converged and reliable parameters.
IV.2 Spectral fitting results
@@
In the upper panels of Figure 2, we compare stellar parameters from our analysis to those calculated with the Gaia XP spectra using the method from Andrae et al. (2023). We find general agreement between the two works. However, there are two notable discrepancies. Below $\sim 5500\,$K, we find a roughly linearly increasing discrepancy in $T_{\rm eff}$, where our estimates are consistently hotter. This is not unexpected given previous work which has also found that stellar parameters determined using different methods and codes tend to become more discrepant with each other at low temperatures (e.g. Hegedűs et al., 2023). We use the spectroscopic $T_{\rm eff}$ measurements throughout for consistency. In Appendix A.4 (Figure 16), we show that our conclusions about which sources have s-process enhancement are not sensitive to this choice.
@@
In addition, our metallicities are systematically lower by 0.2 dex on average than those estimated by Andrae et al. (2023). One factor that may contribute to this is that Andrae et al. (2023) provides the mean metallicity, [M/H], not [Fe/H], meaning that it is not a one-to-one comparison. In particular, old stars deficient in iron ([Fe/H]<1.0<-1.0) are usually more enriched in other metals, such as $\alpha$-elements (e.g. Cayrel et al., 2004; Adibekyan et al., 2012) and carbon (e.g. Li et al., 2022), which would increase the average metal content.
@@
For objects found in the fourth data release of the Galactic Archaeology with HERMES (GALAH) Survey (Buder et al., 2025), we also compare the quoted stellar parameters from this catalog (with quality flags flag_sp = 0 and flag_fe_h = 0) to our results in the lower panels. We once again note an offset in [Fe/H]. We discuss how this offset affects our final barium dwarf fractions in Section V.5.
@@
While these comparisons are helpful for validation, we emphasize that Andrae et al. (2023), GALAH, and our work all use distinct methods to estimate stellar parameters and metallicities (data-driven machine learning, synthetic spectral fitting with neural networks, and excitation/ionization balance, respectively) and different underlying radiative transfer codes, meaning systematic offsets are not unexpected (e.g. Blanco-Cuaresma, 2019).
@@
In Figure 3, we compare regions of the observed spectra for three objects to synthentic spectra generated using our best-fit $T_{\rm eff}$, log(g)(g), and [Fe/H]. We generated the synthetic spectra using the iSpec (Blanco-Cuaresma et al., 2014; Blanco-Cuaresma, 2019), which provides a python interface through which we employed the SPECTRUM radiative transfer code [^4] https://www.appstate.edu/ grayro/spectrum/spectrum.html (Gray and Corbally, 1994). The projected rotational velocity was fixed to the solar value ($v\sin(i)=1.6\,$km/s; e.g. Pavlenko et al. 2012) and the resolution was set to 50,000. We did not vary individual elemental abundances when generating these models. We include these plots for objects observed with different instruments and at a range of metallicities to provide visual confirmation that the best-fit parameters obtained through our method described above are reasonable.
@@
Figure 3: Observed spectra of three targets (gray) compared with synthetic spectra generated using their best-fit stellar parameters (pink; Section IV.1). We select targets covering a wide range of [Fe/H] with similar $T_{\rm eff}$ and log(g)(g), observed with different instruments. On the left, we plot the region around the magnesium triplet, and on the right, we move to redder wavelengths populated with iron lines. Visually, we see good agreement between the observations and the corresponding models, suggesting that our stellar parameters are reasonable. For the most metal-poor star (top row), the C2 Swan band at 51655165\,Å is visible, which suggests carbon enhancement (Section V.4).
IV.3 Deriving abundances
@@
Together with the best-fit stellar parameters determined above, the EWs are inputted into q2, which returns absolute abundances of specified elements inferred by each line. We consider lines with EWs between 20 and 180180\,mÅ. Compared to iron, we set a stricter lower bound to avoid the weakest lines where noise is being fit; while we set a less restrictive upper bound as there are fewer strong lines. Most elements of interest have at least one line with EW below this looser upper limit, and we have confirmed that the inferred abundances from individual lines with EWs above and below 120120\,mÅ are consistent, suggesting that saturation is not a major issue. In Table 1, we list key lines of elements used to determine their abundances which are discussed in Section V. Note that we completely exclude several strong lines (e.g. Ba II $\lambda$4554Å) as these are often saturated, making their EWs uninformative.
For each element, we discard lines which have residuals of more than 0.3 dex from the median. Unless otherwise stated, we use solar abundances from Asplund et al. (2021) to calculate the differential abundances. Where there are multiple lines, the best-fit value is taken to be the median of the abundances inferred from each line and the error is taken to be their standard deviation divided by the square root of the total number of lines. If only one line is measured, we set a fixed error of 0.2 dex. This is a conservative choice, being roughly the 90th percentile of the standard deviations of abundances with multiple lines.
We note that as expected, there is close agreement between [Fe I/H] and [Fe II/H] for most of our systems. However, approximately 17% of them have discrepancies between greater than 0.20.2\,dex. Almost all of these are cases where log(g)(g) was fixed to achieve convergence of stellar parameters (Section IV.1). Since log(g)(g) is determined via ionization balance, this is unsurprising.
@@
V Results
The resulting best-fit stellar parameters and abundances for all 160 objects are summarized in Tables 4 and 5 in Appendix A.5.
V.1 Abundance patterns
@@
Panels (B) to (D) show $\alpha$ element abundances relative to iron, $\alpha$/Fe], against [Fe I/H]. Our measured abundances follow the typical anti-correlation between $\alpha$/Fe] and [Fe/H] observed in the Galactic disk (e.g. Hayden et al., 2015), with a high-$\alpha$ plateau at low [Fe/H] due to rapid enrichment of stars formed at early times by core collapse supernovae (e.g. Burbidge et al., 1957; Tinsley, 1979, 1980; Delgado Mena et al., 2017; Buder et al., 2021). This is one indication that our measured abundances are reasonable.
@@
[…]
In the right panel of Figure 6, we plot [Ba/Y] (i.e. [Ba/Fe] - [Y/Fe]) as a function of the metallicity, [Fe/H]. This is a measure of the relative abundance of heavy (e.g. Ba, La, Ce) to light s-process (e.g. Y, Zr) elements (often denoted as [hs/ls]), which is commonly used to compare observations to predicted yields from AGB nucleosynthetic models (e.g. Lugaro et al., 2003; Allen and Barbuy, 2006). Above [Fe/H]=1.0=-1.0, almost all stars are consistent with negative [Ba/Y] within their observational uncertainties, while the two stars with [Fe/H]<1.0<-1.0 have positive [Ba/Y]. This implies that the yields of the light s-process element, Y, is greater than heavy s-process element, Ba, in AGB donors, unless they are very metal-poor. Although this trend between [Fe/H] and [Ba/Y] is in the expected direction based on previous work, we note that it is weaker than found in most previous work, in particular below [Fe/H]=1.0=-1.0 (e.g. Smiljanic et al., 2007; Cristallo et al., 2016; Cseh et al., 2018; Yang et al., 2024).
@@
We find no clear separation in the abundance patterns for systems in the NCE vs. RCE samples, which suggests that – within our sample, which is restricted to RCE stars with relatively small color excesses, low eccentricity, and companion masses near $0.6\,M_{\odot}$ – they probe the same underlying population. In particular, there are barium dwarfs in the RCE sample, which would not be expected if they were triples containing tight MS+MS binaries. This reflects the fact that the cut on color excess removes some fraction of true WD + MS binaries from the non-class I sample, as discussed in Yamaguchi et al. (2025).
@@
V.2 Dependence on orbital and stellar parameters
In Figure 8, we show the measured [Ba/Fe] against several orbital and stellar parameters: orbital period ($P_{\rm orb}$), photocenter semi-major axis ($a_{0}\times d$, where $a_{0}$ is the angular size in mas and $d$ is the distance in kpc; this is equal to the primary semi-major axis for a dark secondary), eccentricity ($e$), MS star mass ($M_{\rm 1}$), and companion mass ($M_{\rm 2,dark}$). $M_{\rm 1}$ was obtained from Gaia’s binary_masses catalog. $M_{\rm 2,dark}$ is the mass inferred from the astrometric orbit assuming a completely dark companion, which is a reasonable assumption for a cool WD or NS. To estimate the uncertainty, we use Monte Carlo sampling, generating 100 realizations of $M_{\rm 2,dark}$ by drawing from Gaussian distributions of the input parameters ($a_{0}$, $P_{\rm orb}$, $M_{\rm 1}$, and parallax $\varpi$) and taking the standard deviation of these realizations. The histogram under each plot shows the fraction of barium dwarfs (as defined in Section V.1) over the total number of stars in our sample. We selected bins so that the number of systems in each bin is approximately constant. The same plots for [Y/Fe] can be found in Figure 17 in Appendix A.6.
Overall, we see no strong dependence of either [Ba/Fe] or the barium dwarf fraction with orbital period or semi-major axis. There is no strong trend with eccentricity at $e<0.3$, but at higher eccentricities, we find only one barium star (Section V.3). This is likely due to highly eccentric systems hosting NS or high-mass WD secondaries, which originate from more massive progenitors with lower s-process yields (Section VII).
In panels (D) and (E.2), we see that for barium dwarfs, [Ba/Fe] increases with both increasing primary and secondary mass, consistent with the findings of Rekhi et al. (2024, 2025). It is important to point out that these quantities are not independent of each other. More massive primaries require more massive WDs to rule out single MS companions and enter the non-class I sample.
However, the trend with secondary mass eventually falls off for masses above $\sim 0.8\,M_{\odot}$. Moreover, across the entire range of companion masses, there are stars which show no s-process enhancement. This suggests that mass is not the sole determining factor for whether or not stars end up as barium dwarfs. We discuss possible physical explanations of the observed trends and scatter in Section VII.